# SSupplemental References

### v2

@@ 3.5 Comparison of analysis scores
Fig 12. These figures show the comparison charts for the different variants of AHP with the original AHP across the Ganga buffer. We used three map types to highlight the differences between the approaches. Column (a) is Map type 1 (colors spread between overall min max), Column (b) is Map type 2 (colors spread between local min max), and Column (c) is Map type 3 (colors spread between 5-95 percentile). Rows 1-8 are Nested AHP, ANP, 1-N AHP (Worst case), 1-N AHP (Avg case), Fuzzy AHP (Mean), Fuzzy AHP (Case1), Fuzzy AHP (Case2) and Population density, respectively.
3.5.1 Difference between Nested AHP and ANP
Observing Nested AHP and ANP, we saw a similar pattern of differences with AHP, owing to the non-linear distribution of both the results. While the AHP predicted the vulnerability to be high in Kanpur, ANP and Nested AHP suggest that it would be lower. And the vice versa is seen to be true for Prayagraj/Allahabad and Varanasi, where AHP predicted the vulnerability to be lower than the results of ANP and Nested AHP.
3.5.2 Difference between 1-N AHP and AHP
Despite the different visualization approaches, there were no major differences for the vulnerability of 1-N AHP with AHP. This can be explained by the probabilistic layers of acute and chronic error, and the fact that the factors other than acute and chronic are strongly correlated with the AHP variables.
3.5.3 Difference between Fuzzy AHP cases (Mean, Case 1, and Case 2) and AHP
In the Mean Fuzzy AHP Case, we sampled some of the simulations, and most of them had a rank reversal between rainfall and drainage density, causing those layers to reduce the robustness of the results. In Case 1 and Case 2, both approaches were in agreement with each other across the latter half of the Ganga buffer, but we see that in the high-altitude areas they were tending in opposite directions. In Case 1 the AHP suggested a higher vulnerability than Fuzzy AHP, primarily affected by Population Density, and in Case 2, Fuzzy AHP returned a higher value for vulnerability scores, owing to the Annual Rainfall variable.
3.5.4 Difference between Population density and AHP
The differences between the standardized values of AHP and the Population density layers were the highest among all approaches, and there were many areas where there is a non-zero value for this layer. This indicates that AHP is amply influenced by the other parameters and proves the effectiveness of the AHP approaches for creating a good composite variable.
3.5.5 Analysis of Variances
To compare the predictions for multiple variants at once, we calculated the standard deviation between three approaches: AHP, ANP, Nested AHP for each pixel and visualized them for our region of interest (Fig 13). We noticed that the deviations were high for the tier-two cities Prayagraj/Allahabad and Varanasi. Varanasi had the highest vulnerability in terms of AHP, but the other variants of AHP brought more emphasis to Prayagraj/Allahabad as well. There were also some high deviations in the high-altitude regions near Rishikesh - a high-population city in Uttarakhand, suggesting disagreements between the approaches.
Fig 13. Standard deviations of Normalized layers for the approaches AHP, ANP, Nested AHP, where darker values suggest lower standard deviation, and the lighter values suggest high-standard deviations.
3.6 Composite Layer for all Approaches
To create a composite layer to capture the effect of each variant of AHP across the river buffer, we took a normalized base layer for AHP, and overlaid the outliers (>=2 standard deviations) for each of the difference layers for Mean Fuzzy AHP, Nested AHP, and ANP (Fig 14). To understand the distribution of the AHP base layer, and the difference layers for Mean Fuzzy AHP, Nested AHP, and ANP, we also created the histograms of each layer (Fig 15). In the composite layer, the primary colors (red, blue, yellow) were used to depict the outliers for the individual layers, and secondary colors (orange, green, purple) were used when the effects of two AHP variants were combined. Given that the outliers could be both positive and negative, we also created separate maps for the two. These composite layers can be used as a way to depict a single variable for river Ganga’s vulnerability with the effect of all AHP variants. Through these composite layers, we are illustrating the effect of AHP relative to the unique effects of the variants wherever they disagree with AHP.
We also selected four important subsections of the river buffer where the composite variable was effective in capturing the effects of each AHP variation (AHP, Nested AHP, and ANP). There are some small areas near the river body which had all three outliers. We can see the effect of the combination of two variants in Fig 16 subsection C, where there were a lot of outliers for both Nested AHP + ANP (violet) and Nested AHP + Mean Fuzzy AHP (green) variants.
Fig 14. The composite layer (a) that can be used to describe the vulnerability of the river Ganga. We assign different colors for the outliers of each variant, and to showcase the positive and negative outliers, we split them into two charts (b) and (c)
Fig 15. Here, (a) shows the Normalized AHP layer with its standard deviation, while the rest of the figures show the difference between Normalized AHP and its corresponding AHP Variants: ANP, Mean Fuzzy AHP, and Nested AHP. (b) is Normalized AHP minus Normalized ANP. (c) is Normalized AHP minus Normalized Mean Fuzzy AHP. (d) is Normalized AHP minus Normalized Nested AHP
Fig 16. Subsection analysis of the composite variable for four subsections A, B, C, D, with different panels numbered 1 (all outliers), 2 (positive outliers), and 3 (negative outliers). To cover a range of distributions, we considered two rural (A, D), and two urban (B, C) subsections. Out of these, Subsection B has the city Kanpur in them, while subsection C is near Prayagraj/Allahabad and Varanasi, highly populated areas along the Ganga.
4 Discussion
In this paper, we aimed to assess the pollution vulnerability levels of the river Ganga using remote sensing datasets. We lacked ground truth data for pollution, making supervised learning algorithms unfeasible, and making it difficult to validate the metric. Gaining confidence around our results was critical, as decisions regarding the river’s ecological health depend on a reliable assessment of vulnerability levels. In order to generate a spatial numeric dataset reflecting the pollution vulnerability of the river Ganga, we utilized the widely used AHP decision making method.
Despite the advantages of AHP, it has multiple shortcomings - one being its reliance on expert opinion which might be biased. So we explored existing alternative methods, such as ANP and Fuzzy AHP, and we designed novel variants of AHP, such as Nested AHP, 1-N AHP, and a hybrid Fuzzy 1-N AHP. Each of these approaches offered unique advantages over the original AHP approach, and contributed towards the robustness of our analysis by addressing different shortcomings of the AHP method.
For instance, AHP functions under the assumption that the criteria are independent of each other (5.2.8), which is fairly uncommon to achieve in real world scenarios, especially in ecological settings. For this reason, we used ANP to capture the criteria dependencies, and the results of ANP were accommodated into our proposed solution. Another shortcoming being addressed by our approach is that we tested the uncertainty in the decision space (5.2.24) by checking the likelihood of rank reversals using Fuzzy AHP and Fuzzy 1-N AHP, thus ensuring that our solution is robust to uncertainties in the decision space. Within Fuzzy AHP, we introduced another novel methodology which involved using Monte Carlo simulations to get randomized values of fuzziness instead of relying on the domain experts. This way we managed to quickly run a variety of scenarios of AHP while excluding the need of domain experts. Lastly, the “external” influence on decision (5.2.26) is accommodated by expanding to the breadth of all possible factors that might influence the decision space using 1-N AHP.
Although this study addressed many of AHP’s known shortcomings, there were some limitations that were not addressed. Of note, the need for multiple technical experts to address the broad range of factors when assigning Saaty rankings is a valid concern. While we are an interdisciplinary group, we are not subject matter experts across all of the fields represented by the AHP input factors. To address this issue, an overall quality metric of the Saaty rankings may mitigate this issue based on the level of expertise and available data (e.g., a quality metric might consist of four levels, ranging from novice opinion to the consensus of subject matter experts). Furthermore, lack of certainty about the importance of alternative criteria and biases during ranking may be addressed by running different scenarios, as long as the decision space is fundamentally partitioned based on an overarching difference due to a given circumstance, and robust data and guidance is provided when rankings are calculated. In addition, vulnerability scores were dependent on remote sensing data resolution (e.g., population density is reported at 1 km2 scale) and were context sensitive such that decision spaces may vary depending on the influence, conditions, and/or regulatory as well as technical mitigation potential in different biomes, nations, and jurisdictions to alleviate sources of vulnerability. Finally, dependencies and correlations between or among factors may not be well understood. While this is true, the process is iterative such that the models can be refined as additional data are gathered, factor interactions are clarified, and scenarios are tested.
Additional data layers such as GIS-based applications that track impacts may be integrated into the data process to support and hone vulnerability assessments and thus AHP factor weights as well as improve output resolution and validation. For instance, acute factors may be known unknowns (e.g., mining locations) and unknown/ill defined unknowns (e.g., location and breadth of endangered species habitat). Similarly, potential widespread impacts such as loss of biodiversity and fragmentation of biodiversity richness are chronic factors that were excluded from our AHP vulnerability decision space but may be included in future efforts. Indeed, the approach of sand mining and waste disposal tracked by Bayazidy et al. (2024) may be used to identify specific instances of these and other acute factors that require precise and timely identification. In addition, coupling biodiversity and the conservation imperative analysis of Dinerstein et al. (2024) with our vulnerability assessment can further prioritize areas for conservation that were scored as low biodiversity and extremely low vulnerability. Moreover, emerging factors may be added as data are collected and the breadth and depth of the decision space grow. For instance, the environmental spread of antibiotic resistance in freshwater is tracked with molecular genetic techniques (Yang et al., 2018; Loi, 2023), and coupling these data with their spatial locations can add antibiotic resistance as a vulnerability factor. Finally, high-resolution data from UAV such as that produced by Tripathi et al. (2024) during their effort to combine remote sensing with drone data while cataloging riparian vegetation along the Ganga can augment and further refine LULC information to better refine vulnerability scores.
Ultimately, networks of models may be brought together to enrich landscape digital twins for design, decision making, and monitoring of real world features of landscapes. These twins become digital representations of natural and built features in landscapes. For river systems, digital twins may consist of riverscapes, the concept that integrates longitudinal characteristics of rivers with their land features (Zhou et al., 2014). Indeed, the basis for such twins already exists in regions prone to flooding. For instance, Nested AHP was used to predict five categories of flood susceptibility with nine factors based on remote sensing data in addition to soil texture, geomorphology, and geology (Das and Gupta, 2021). Furthermore, AHP with nesting of four factors (land use, percent green space, per capita sewer length, and slope) was used to produce a flood vulnerability index in a section of Hanoi, Vietnam (Liyanage and Yamada, 2017), and a flood susceptibility map for all of Hanoi was made with a 9-factor AHP that was ranked by subject matter experts (Nguyen et al., 2024). AHP models that include cultural, socioeconomic and environmental criteria provide for wider assessments of the human condition. Recently, 20 factors representing these criteria as well as security and service functions were placed in a hierarchical structure and assessed through a survey of subject matter experts to realize a quality of life score (Withanage et al., 2024). Of note, they detected the inverse of our vulnerability scores in that quality of life tended to decrease away from city centers and the services they provide whereas cities consistently were sources of high vulnerability to the river Ganga. Thus, the perspective of the modeling scenario must be understood and integrated into the larger decision space of balance between human endeavors and nature. A richness of alternative perspectives can drive sustainable development so that factors that improve quality of life can be implemented with policies that recognize and mitigate their accompanying sources of vulnerabilities.
5 Conclusion
In this study, we assessed the vulnerability of the river Ganga to pollution along a 20 km wide corridor covering 26,609.4 km2 of riverscape. We used AHP to perform a dimensionality reduction of six remote sensing datasets to create a single variable for river vulnerability. To mitigate structural limitations to AHP, our approach brought together a suite of methods and produced a set of comparative metrics to identify and refine vulnerability scores. Of note, urban landscape features with extremely high vulnerability scores and those at the interface with vulnerable areas were identified. These findings provide the basis to rational prioritization of pollution mitigation and a list of locations for future rejuvenation. Moreover, sites currently with low vulnerability scores may be targeted for conservation or sustainable development practices to prevent their degradation. Our vulnerability score, together with other metrics such as a quality of life score (Withanage et al., 2024), can enrich digital twins beyond simplistic animations and reveal their underlying environmental, social, and economic characteristics.
6 Declaration of competing interest
The authors declare that they have no conflicting interests.
7 Acknowledgements
The authors thank Markley Boyer and Sarah Bergmann for thoughtful discussions throughout the project. We also thank Aaron Hirsh for organizing Collaborative Earth.
[…]
SUPPLEMENTAL INFORMATION
Table S1: Shortcomings of AHP analysis rubric.
5.2.1 The Pair-Wise Method and Its Application in AHP N
5.2.2 The Pair-Wise Method in AHP Constructs Artificial Relationships Y Test the robustness of the analysis, and define its uncertainty as well as consistency. Fuzzy AHP
5.2.3 Criteria Preferences Must Consider Alternatives N
5.2.4 The Ambiguity of Pair-Wise Comparisons in AHP N
5.2.5 Modelling Scenarios N .
5.2.6 Lack of Rational Answers from DMs N
5.2.7 AHP Incapacity to Solve Complex Problems Y Complex problems have trade-offs of varying importance and levels of certainty 1-N AHP, ANP, Fuzzy AHP
5.2.8 Criteria Independency Y A testing of marginal effects is warranted to understand the strengths of interactions. 1-N AHP, ANP
5.2.9 Quantifying Preferences N
5.2.10 Quantitative Data N
5.2.11 The DM Should Be a Multiple Technical Expert Y It’s a valid concern and surveys of experts to find a consensus (and range) of expert assessments.
5.2.12 The AHP Tries to Subordinate Reality to a Theory Y Using different decision matrices for different categories of decision (e.g., cost, siting, restoration category) is reasonable. 1-N AHP
[…]
5.2.14 AHP’s Fundamental Scale Has Limits 1 and 9 N
5.2.15 A Preference and the Peculiar Meaning of Its Inverse Value N
5.2.16 Determination of Criteria Importance Must Contemplate the Alternatives Y Different scenarios should be run. If the decision space is fundamentally partitioned based on an overarching difference due to a given circumstance.
Table S2: Shortcomings of AHP analysis rubric (continued).
5.2.19 Normalization Y 1-N AHP is a test of this. 1-N AHP
5.2.20 The Use of the Eigenvector Method N
5.2.27 Not all problems have a hierarchy N
5.2.28 Criteria limiting other criteria N
5.2.29 Problem structuring N
Fig S1. Panels (a) through (e) are box and whisker plots for the AHP factors Population Density (a), Annual Rainfall (b), Drainage Density (c), Slope (d), and Land Surface Temperature (e) for the length of the Ganga contained in the area of analysis. Data distribution was tested and results are given for each factor label. Whiskers show the range of values, the box "shoulders" are the inter-quartile range, and the median is shown as a band.
Fig S2. Panels (a) through (l) are the histograms and log-histograms to assess the distributions of the AHP factors along the length of the Ganga contained in the area of analysis. Panels (a), (c), (e), (g), (i) represent the histograms for Population Density, Annual Rainfall, Drainage Density, Slope, and Land Surface Temperature, respectively, while panels (b), (d), (f), (h), (j) respectively represent the log-histograms for the same. Panels (a) through (j) are binned using deciles, and coloured green to red for low to high vulnerability. (k) and (l) represent LULC and are color coded to their respective Land Use classes.
Table S3: Data Sources
Name Description Locationa References
Buffer Ganga river This study This study
[…]
aData identified as from this study is available at Drive Link
AHP pairwise comparisons
[…]
Table S4: AHP Pairwise and Normalized Comparison Matrixa
a.) AHP pairwise comparison matrix
PD LULC RAIN DD SLOPE LST
PD 1 2 4 5 7 9
LULC 0.5 1 3 4 5 7
RAIN 0.25 0.33 1 2 4 5
DD 0.2 0.25 0.5 1 3 5
SLOPE 0.14 0.2 0.25 0.33 1 4
LST 0.11 0.14 0.2 0.2 0.25 1
Total 2.2 3.92 8.95 12.53 20.25 31
b.) Normalized AHP pairwise comparison matrix PD LULC RAIN DD SLOPE LST Mean normalized factor weight PD 0.45 0.51 0.45 0.4 0.35 0.29 0.41 LULC 0.23 0.26 0.34 0.32 0.25 0.23 0.27 RAIN 0.11 0.08 0.11 0.16 0.2 0.16 0.14 DD 0.09 0.06 0.06 0.08 0.15 0.16 0.1 SLOPE 0.06 0.05 0.03 0.03 0.05 0.13 0.06 LST 0.05 0.04 0.02 0.02 0.01 0.03 0.03 aPairwise comparison matrix for AHP factors, their normalized weight matrix, and overall mean normalized weights which were used to determine a vulnerability score per pixel. PD, population density; LULC, land use and land change; RAIN, rainfall; DD, drainage density; LST, land surface temperature.
Natural breaks and AHP factor classes
[…]
Nested AHP pairwise comparisons
[…]
Table S5: Nested AHP: Factor Vulnerability Rankings for all Factors
LULC Built-up Cropland Barren/ Sparse Vegetation Trees/ Grass/ Shrubs Water/ Wetland
Built-up 1 3 4 5 7
Cropland 0.33 1 1 2 4
Barren/Sparse Vegetation 0.25 1 1 1 3
Trees/ Grass/ Shrubs 0.2 0.5 1 1 2
Water / Wetland 0.14 0.25 0.33 0.5 1
Total 1.93 5.75 7.33 10 17
Rainfall (mm/y) >1410 1189-1410 1019-1189 868-1019 <868
>1410 1 3 4 5 5
1189-1410 0.33 1 1 2 2
1019-1189 0.25 1 1 1 1
868-1019 0.2 0.5 1 1 1
<868 0.2 0.5 1 1 1
Total 1.98 6 8 10 10
Drainage density (km/km2) >4 2.6-4 1.5-2.6 0.6-1.5 <0.6
>4 1 6 8 9 9
2.6-4 0.17 1 2 3 3
1.5-2.6 0.13 0.5 1 1 1
0.6-1.5 0.11 0.33 1 1 1
<0.6 0.11 0.33 1 1 1
Total 1.51 8.17 13 15 15
Slope (Angle %) >35 24-35 45650 45394 <4
>35 1 1.5 3 6 9
24-35 0.67 1 2 5 8
45650 0.33 0.67 1 4 7
45394 0.17 0.22 0.25 1 3
<4 0.11 0.13 0.14 0.33 1
Total 2.28 3.52 5.89 16 28
Temperature (℃) >37 34-37 30-34 25-30 <25
>37 1 4 5 6 8
34-37 0.25 1 1 2 4
30-34 0.2 1 1 1 3
25-30 0.17 0.5 1 1 2
<25 0.13 0.25 0.33 0.5 1
Total 1.74 6.75 8.33 11 18
ANP supermatrix and inner dependence (W22 and W33) pairwise comparisons
[…]
Fig S3. Inner dependencies between ANP criteria (AHP factors) are shown as ovals. Arrows about the ovals indicate influence by a criterion on other criteria to its right and arrows below the ovals indicate influence by a criterion on other criteria to its left.
Fig S4. These figures show the ANP’s vulnerability score with the original AHP submatrix; the lower values pertained to the least vulnerable areas while the ones with higher scores were the most vulnerable. Map of exact values per pixel location (a) and histogram of binned ranges of values (b).
[…]

### v3

@@ 1-N AHP calculation
Original AHP & AHP Variants Comparison Chart These figures show the comparison charts for the different variants of AHP with the original AHP across the Ganga buffer. We used three map types to highlight the distribution of differences between the approaches. Column (a) is Map type 1 (colors spread between overall min max), Column (b) is Map type 2 (colors spread between local min max), and Column (c) is Map type 3 (colors spread between 5-95 percentile). Rows 1-8 are Nested AHP, ANP, 1-N AHP (Worst case), 1-N AHP (Avg case), Fuzzy AHP (Mean), Fuzzy AHP (Case1), Fuzzy AHP (Case2) and Population density, respectively.
S7 Text
Difference between Fuzzy 1-N AHP, Fuzzy AHP and 1-N AHP
While Fuzzy 1-N AHP gave us some insights on the combined effects of Fuzzy AHP and 1-N AHP, to test whether it is greater than the sum of its parts, which would warrant an additional comparison on its value addition to AHP, we ran a two-way ANOVA test with the parameters quintile and model type (Fuzzy 1-N AHP, Mean Fuzzy AHP, 1-N AHP Worst case) (S7 Table.). While the p-values for quintiles, and the combined effect of quintile and model type on the vulnerability scores were <0.0001, the p-value for models only was 0.9270. When we performed a Tukey post-hoc test for each quintile, we found that (other than in quintile 1) there was no significant difference between Fuzzy 1-N AHP and Fuzzy AHP. Quintiles 1, 4, and 5 of the Fuzzy 1-N AHP and 1-N AHP comparison were significantly different which reinforced that the edges of the distribution were susceptible to variability caused by uncertainty in model factors as was seen in the 1-N AHP and AHP comparison. Together, these results suggested that the individual analyses of Fuzzy AHP and 1-N AHP can represent the effect of Fuzzy 1-N AHP in comparison with classic AHP. Hence we did not further include Fuzzy 1-N AHP in the comparison analysis or S5 Fig..
S5 Table.
Two-way ANOVA and Tukey Post-Hoc test for AHP, ANP and Nested AHP Sum of Squares DF F statistic p-value Quintile 18554.85 4 6539.64 <0.0001<0.0001 Model 0.51 2 0.36 0.6977702 Quintile + Model 1065.47 8 187.76 <0.0001<0.0001 Residual 10629.19 14985
Quintile 1
Group 1 Group 2 Mean Diff p-adj Lower Upper Reject
AHP ANP 0.458 <0.0001<0.0001 0.3748 0.5412 TRUE
AHP Nested AHP 0.781 <0.0001<0.0001 0.6978 0.8642 TRUE
ANP Nested AHP 0.323 <0.0001<0.0001 0.2398 0.4062 TRUE
Quintile 2
Group 1 Group 2 Mean Diff p-adj Lower Upper Reject
AHP ANP 0.506 <0.0001<0.0001 0.3949 0.6171 TRUE
AHP Nested AHP 0.748 <0.0001<0.0001 0.6369 0.8591 TRUE
ANP Nested AHP 0.242 <0.0001<0.0001 0.1309 0.3531 TRUE
Quintile 3
Group 1 Group 2 Mean Diff p-adj Lower Upper Reject
AHP ANP -0.161 0.0005 -0.2616 -0.0604 TRUE
AHP Nested AHP -0.427 <0.0001<0.0001 -0.5276 -0.3264 TRUE
ANP Nested AHP -0.266 <0.0001<0.0001 -0.3666 -0.1654 TRUE
Quintile 4
Group 1 Group 2 Mean Diff p-adj Lower Upper Reject
AHP ANP -0.471 <0.0001<0.0001 -0.5531 -0.3889 TRUE
AHP Nested AHP -0.668 <0.0001<0.0001 -0.7501 -0.5859 TRUE
ANP Nested AHP -0.197 <0.0001<0.0001 -0.2791 -0.1149 TRUE
Quintile 5
Group 1 Group 2 Mean Diff p-adj Lower Upper Reject
AHP ANP -0.385 <0.0001<0.0001 -0.4387 -0.3313 TRUE
AHP Nested AHP -0.502 <0.0001<0.0001 -0.5557 -0.4483 TRUE
ANP Nested AHP -0.117 <0.0001<0.0001 -0.1707 -0.0633 TRUE
S6 Table.
Two-way ANOVA and Tukey Post-Hoc test for AHP, Mean Fuzzy AHP, Fuzzy AHP Case 1 and Fuzzy AHP Case 2
Sum of Squares DF F statistic p-value
Quintile 31324.9 4 18828.71 <0.0001<0.0001
Model 7.94 3 6.36 <0.0001<0.0001
Quintile + Model 534.84 12 107.16 <0.0001<0.0001
Residual 8310.07 19980
Quintile 1
Group 1 Group 2 Mean Diff p-adj Lower Upper Reject
AHP Mean Fuzzy AHP 0.037 0.3115 -0.0182 0.0922 FALSE
AHP Fuzzy AHP Case 1 0.28 <0.0001<0.0001 0.2248 0.3352 TRUE
AHP Fuzzy AHP Case 2 0.549 <0.0001<0.0001 0.4938 0.6042 TRUE
Mean Fuzzy AHP Fuzzy AHP Case 1 0.243 <0.0001<0.0001 0.1878 0.2982 TRUE
Mean Fuzzy AHP Fuzzy AHP Case 2 0.512 <0.0001<0.0001 0.4568 0.5672 TRUE
Fuzzy AHP Case 1 Fuzzy AHP Case 2 0.269 <0.0001<0.0001 0.2138 0.3242 TRUE
Quintile 2
Group 1 Group 2 Mean Diff p-adj Lower Upper Reject
AHP Mean Fuzzy AHP 0.095 0.0275 0.0073 0.1827 TRUE
AHP Fuzzy AHP Case 1 0.04 0.6441 -0.0477 0.1277 FALSE
AHP Fuzzy AHP Case 2 0.304 <0.0001<0.0001 0.2163 0.3917 TRUE
Mean Fuzzy AHP Fuzzy AHP Case 1 -0.055 0.3715 -0.1427 0.0327 FALSE
Mean Fuzzy AHP Fuzzy AHP Case 2 0.209 <0.0001<0.0001 0.1213 0.2967 TRUE
Fuzzy AHP Case 1 Fuzzy AHP Case 2 0.264 <0.0001<0.0001 0.1763 0.3517 TRUE
Quintile 3
Group 1 Group 2 Mean Diff p-adj Lower Upper Reject
AHP Mean Fuzzy AHP 0.053 0.3898 -0.0332 0.1392 FALSE
AHP Fuzzy AHP Case 1 0.045 0.5361 -0.0412 0.1312 FALSE
AHP Fuzzy AHP Case 2 -0.117 0.0028 -0.2032 -0.0308 TRUE
Mean Fuzzy AHP Fuzzy AHP Case 1 -0.008 0.9952 -0.0942 0.0782 FALSE
Mean Fuzzy AHP Fuzzy AHP Case 2 -0.17 <0.0001<0.0001 -0.2562 -0.0838 TRUE
Fuzzy AHP Case 1 Fuzzy AHP Case 2 -0.162 <0.0001<0.0001 -0.2482 -0.0758 TRUE
Quintile 4
Group 1 Group 2 Mean Diff p-adj Lower Upper Reject
AHP Mean Fuzzy AHP -0.129 0.0002 -0.2097 -0.0483 TRUE
AHP Fuzzy AHP Case 1 -0.209 <0.0001<0.0001 -0.2897 -0.1283 TRUE
AHP Fuzzy AHP Case 2 -0.541 <0.0001<0.0001 -0.6217 -0.4603 TRUE
Mean Fuzzy AHP Fuzzy AHP Case 1 -0.08 0.0529 -0.1607 0.0007 FALSE
Mean Fuzzy AHP Fuzzy AHP Case 2 -0.412 <0.0001<0.0001 -0.4927 -0.3313 TRUE
Fuzzy AHP Case 1 Fuzzy AHP Case 2 -0.332 <0.0001<0.0001 -0.4127 -0.2513 TRUE
Quintile 5
Group 1 Group 2 Mean Diff p-adj Lower Upper Reject
AHP Mean Fuzzy AHP -0.104 <0.0001<0.0001 -0.157 -0.051 TRUE
AHP Fuzzy AHP Case 1 -0.227 <0.0001<0.0001 -0.28 -0.174 TRUE
AHP Fuzzy AHP Case 2 -0.457 <0.0001<0.0001 -0.51 -0.404 TRUE
Mean Fuzzy AHP Fuzzy AHP Case 1 -0.123 <0.0001<0.0001 -0.176 -0.07 TRUE
Mean Fuzzy AHP Fuzzy AHP Case 2 -0.353 <0.0001<0.0001 -0.406 -0.3 TRUE
Fuzzy AHP Case 1 Fuzzy AHP Case 2 -0.23 <0.0001<0.0001 -0.283 -0.177 TRUE
S7 Table.
Two-way ANOVA and Tukey Post-Hoc test for Fuzzy 1-N AHP, Fuzzy AHP and 1-N AHP Sum of Squares DF F statistic p-value Quintile 24588.76 4 17336.98 <0.0001<0.0001 Model 0.05 2 0.08 0.9270276 Quintile + Model 109.21 8 38.5 <0.0001<0.0001 Residual 5313.25 14985
Quintile 1
Group 1 Group 2 Mean Diff p-adj Lower Upper Reject
Fuzzy 1-N AHP Fuzzy AHP -0.096 <0.0001<0.0001 -0.1433 -0.0487 TRUE
Fuzzy 1-N AHP 1-N AHP 0.27 <0.0001<0.0001 0.2227 0.3173 TRUE
Fuzzy AHP 1-N AHP 0.366 <0.0001<0.0001 0.3187 0.4133 TRUE
Quintile 2
Group 1 Group 2 Mean Diff p-adj Lower Upper Reject
Fuzzy 1-N AHP Fuzzy AHP 0.053 0.1918 -0.0186 0.1246 FALSE
Fuzzy 1-N AHP 1-N AHP 0.055 0.169 -0.0166 0.1266 FALSE
Fuzzy AHP 1-N AHP 0.002 0.9976 -0.0696 0.0736 FALSE
Quintile 3
Group 1 Group 2 Mean Diff p-adj Lower Upper Reject
Fuzzy 1-N AHP Fuzzy AHP 0.072 0.0682 -0.0041 0.1481 FALSE
Fuzzy 1-N AHP 1-N AHP -0.007 0.9747 -0.0831 0.0691 FALSE
Fuzzy AHP 1-N AHP -0.079 0.0397 -0.1551 -0.0029 TRUE
Quintile 4
Group 1 Group 2 Mean Diff p-adj Lower Upper Reject
Fuzzy 1-N AHP Fuzzy AHP -0.001 0.9993 -0.0681 0.0661 FALSE
Fuzzy 1-N AHP 1-N AHP -0.099 0.0016 -0.1661 -0.0319 TRUE
Fuzzy AHP 1-N AHP -0.098 0.0018 -0.1651 -0.0309 TRUE
Quintile 5
Group 1 Group 2 Mean Diff p-adj Lower Upper Reject
Fuzzy 1-N AHP Fuzzy AHP -0.009 0.8753 -0.0519 0.0339 FALSE
Fuzzy 1-N AHP 1-N AHP -0.198 <0.0001<0.0001 -0.2409 -0.1551 TRUE
Fuzzy AHP 1-N AHP -0.189 <0.0001<0.0001 -0.2319 -0.1461 TRUE